OpenStudy (anonymous):

HELP PLEASE! Two numbers differ by 57. Their sum is 185. Find the numbers. Choices for the smaller number: 54, 64, 74 (STEP BY STEP TO HELP BETTER UNDERTSAND)

OpenStudy (anonymous):

it's like solving a system , let's take x and y those two numbers, now you express mathematically what you said about their difference and sum can you do that ?

OpenStudy (anonymous):

i dont think so...?

OpenStudy (anonymous):

x and y differ by 57 their sum = 185

OpenStudy (anonymous):

let's express the difference so you said that the two numbers which means x and y diffr by 57 i.e. x-y=57 agree ??

OpenStudy (anonymous):

AGREE

OpenStudy (anonymous):

what about the sum can you figure it out ?

OpenStudy (anonymous):

I HAVE NO IDEA :/

OpenStudy (anonymous):

the sum is expressed by + so thier sum is 185 which means x+y=185 and now is it clear enough ?

OpenStudy (anonymous):

SO WE HAVE TWO EQUATIONS RIGHT?

OpenStudy (anonymous):

right and more that that you have a system x+y=185 x-y=57 do you know how to solve it ?

OpenStudy (anonymous):

there is two ways the first is by replacing a variable by its expression in one of the equations and the second way is to add the both sides of the equations to each other what do you prefer ?

OpenStudy (anonymous):

 IS THIS RIGHT??